There are 258 people, adults pay $8 and the children pay $5 to go to a movie theater. The total sum of money the movie theater made that night from selling tickets was $1431. How many children were there.
Answer by stanbon(75887)

There are 258 people, adults pay $8 and the children pay $5 to go to a movie theater. The total sum of money the movie theater made that night from selling tickets was $1431. How many children were there.
-----------------
Let number of adult tickets sold be "x"; Value of these is 8x dollars
Number of child ticks sold is 258-x; Value of these is 5(258-x)=1290-5x dollars
-----------
EQUATION:
value + value = 1431 dollars
8x + 1290-5x = 1431
3x=141
x=47 (number of adult tickets sold)
258-x=211 (number of child tickets sold)
Cheers,
Stan H.